In an infantile paralysis epidemic 700 people contracted the disease. 400 received no serum treatment and of them 80 became paralyzed. Rest of the people received the serum treatment and among those 70 became paralyzed. can you tell the proportion of people receive and do not receive the serum treatment differ?

Need the solution by 2 hours {assignment time duration}


1
Expert's answer
2021-11-12T15:43:58-0500

total number of people = 700


people that received no serum treatment = 400

people that received no serum and became paralyzed = 80

fraction of people that became paralyzed after receiving no treatment = 80/400 = 1/5


people that received serum treatment = 300

people that received serum and became paralyzed = 70

fraction of people that became paralyzed after receiving the treatment = 70/300 = 7/30


ratio of people that became paralyzed after receiving the treatment to people that became paralyzed after receiving no treatment "= 7\/30 : 1\/5 = 35 : 30"

"= 7: 6"